OpenStudy (aleah54):

FAN AND MEDAL Which of the following functions has the largest value when x = 5? p(x) = 2x2 + 4x + 11 h(x) = 4x s(x) = 10x A.p(x) B.h(x) C.s(x) D.All the functions are equal at x = 5 I think its b

OpenStudy (aleah54):

@Atsie @Azionne @alexandrea.lane @atianaaaa @Alessea101 @abdel008 @Aveline @sleepyjess @ShadowLegendX @samanthagreer @shaik0124 @Serenity74 @ShadowCain @DarkBlueChocobo @DivineSolar @Darby.Allen @Dman8592 @Cuanchi

OpenStudy (shaik0124):

just substitute 5 in the place of x and solve u will find the answer

OpenStudy (shaik0124):

for example k(x)=25x; when x=25 the value of k(x) is k(5)=25*25=625 so k(x)=6625

OpenStudy (shaik0124):

sorry k(x)=625

OpenStudy (aleah54):

So its d

OpenStudy (shaik0124):

no

OpenStudy (shaik0124):

whats the value of p x

OpenStudy (aleah54):

2x2 + 4x + 11

OpenStudy (shaik0124):

calculate now

OpenStudy (aleah54):

ok would it be A

OpenStudy (anonymous):

2x2 + 4(x) +11 4+ 4(5) + 11 = 24 + 11 = 35 4(x) 4(5) = 20 10(x) 10(5) = 50 So the answer is c

OpenStudy (anonymous):

Unless you mean 2 X 2 and not 2 times 2 So 2(5)2 +4(5) + 11 20 + 20 + 11 = 51 So then the answer would be A